OpenStudy (anonymous):

solve 9^2x=66 round to the nearest ten-thousandth no its not 0.9534! HELP MEDALS GIVEN

OpenStudy (jdoe0001):

\(\large \bf 9^{2x}=66 ?\)

OpenStudy (anonymous):

yes ( :

OpenStudy (freckles):

who told you it wasn't 0.9534?

OpenStudy (anonymous):

well i have a quiz and i keep putting that answer in and its saying incorrect i swear i thought it was right idk know what to do??

OpenStudy (freckles):

and the equation @jdoe0001 wrote is no different right?

OpenStudy (jdoe0001):

I assume this is covering logarithms?

OpenStudy (anonymous):

yes its covering logarithms and the equation he put is the correct way to write it

OpenStudy (jdoe0001):

k

OpenStudy (freckles):

your answer looks good you did round it to the nearest ten-thousandths

OpenStudy (anonymous):

i keep putting it in it says oits not correct and yeah i did but i dont know???

OpenStudy (anonymous):

okay nevermind i just put it in and your right~!!!!

OpenStudy (freckles):

I don't suppose you can screenshot it and show us?

OpenStudy (freckles):

oh ok lol

OpenStudy (jdoe0001):

hmmm I do get the same 0.95339702305389887954 rounded to nearest thousands, I gather it'd be 0.9534

OpenStudy (freckles):

you mean ten thousandths?

OpenStudy (anonymous):

can you solve this one 14^x+1=36 thats the last one

OpenStudy (jdoe0001):

hmm maybe... it should be yes... ten thousands

OpenStudy (jdoe0001):

\(\bf 14^x+1=36 \implies 14^x=35\implies log(14^x)=log(35) \\ \quad \\ xlog(14)=log(35)\implies x=\cfrac{log(35)}{log(14)} \)

OpenStudy (freckles):

\[14^{x+1}=36 \\ \text{ or } 14^x +1=36 \] you probably mean the latter I just want to make sure :p

OpenStudy (jdoe0001):

hmmm good one there freckles

OpenStudy (anonymous):

yes i mean the second one lol ( :

OpenStudy (freckles):

cool stuff joe done that above then I assume you have to round or something so whip out the calculator if you need

OpenStudy (anonymous):

so how do i get the decimal i got 1.3472 but its not right : (

OpenStudy (freckles):

what you rounding to? ten-thousandths?

OpenStudy (anonymous):

yes im rounding to the ten-thousandths

OpenStudy (freckles):

1.3472 is correct then

OpenStudy (anonymous):

its saying its wrong??? : (

OpenStudy (freckles):

I don't know your thingy seems weird. You said you typed in that earlier answer and it said it was wrong. Then you typed it in again and it said it was right. but yeah that is the answer

OpenStudy (anonymous):

yeah its like this box on my screen we put the answer in the box then send it it says its incorrect but your right it has to be the quiz thing. thank you so much for the help and verifying that my answers was correct ( :

OpenStudy (freckles):

np